OpenStudy (anonymous):

square root of 3x-4 plus square root of 2+x is equal to 2. Solve for x. I came up with x=-2/5 Is this right?

OpenStudy (anonymous):

Plug the answer back in. Does sqrt(3*-2/5) + sqrt(2 - 2 /5) = 2?

OpenStudy (anonymous):

I mean sqrt(3 * - 2/5 - 4) = sqrt(2 - 2/5) = 2

OpenStudy (anonymous):

Oh, yeah, didn't even think of that! thanks! I'll try it.

OpenStudy (anonymous):

nope doesn't look like it can happen without complex numbers, I'll kep trying.

OpenStudy (anonymous):

\[x=7-4 \sqrt{2}\]